OpenStudy (anonymous):

-SOLVE FOR L & W USING SUBSTITUTION & ELIMINATION METHOD: L×W=1200 π×(W/2π)2×L=600

OpenStudy (pfenn1):

do you understand what they mean by substitution and elimination method?

OpenStudy (pfenn1):

rearrange the first equation to solve for L in terms of W. L=1200/W and plug this expression into the second equation for L and solve

OpenStudy (anonymous):

assume the second one is \[\pi \times (\frac{W}{2\pi})^2\times L=600\] or \[\frac{W^2}{4\pi}\times L=600\] right?

OpenStudy (anonymous):

is it a square, or a two?

OpenStudy (anonymous):

2 pie and then it is squared

OpenStudy (anonymous):

ok so replacing \(L\) by \(\frac{1200}{W}\) you get \[\frac{W^2}{4\pi}\times \frac{1200}{W}=600\]

OpenStudy (anonymous):

cancel a W and solve the linear equation

OpenStudy (anonymous):

you got it from here?

OpenStudy (anonymous):

no :( i suck at this one & its a project & i want to pass so i really need help for this one problem please

OpenStudy (anonymous):

ok we start with this line \[\frac{W^2}{4\pi}\times \frac{1200}{W}=600\] cancel and get \[\frac{W}{4\pi}\times 1200=600\] cancel some more and get \[300\times \frac{W}{\pi}=600\] multiply by \(\pi\) and divide by 300 to get \[W=2\pi\]

OpenStudy (anonymous):

theres suppose to be a pie sign in front of w/4pie

OpenStudy (anonymous):

we canceled that one with one of the ones in the denominator

OpenStudy (anonymous):

then how did u get 300?